# Robust covariance estimation and explainable outlier detection for matrix-valued data

### v1

@@ 5.1 Shapley values for multivariate data
  • Monotonicity: Let $\bm{\mu},\tilde{\bm{\mu}}\in\mathbb{R}^{p}$ be two vectors and $\bm{\Sigma},\tilde{\bm{\Sigma}}\in\operatorname{PSD}(p)$ be two matrices. If
@@
In words, the coordinate $\phi_{k}(x)$ of the Shapley value is the average marginal contribution of the $k$-th variable to the squared Mahlanobis distance and is obtained by averaging over all marginal outlyingness contributions $\Delta_{k}\operatorname{MD}^{2}(\hat{\bm{x}}^{S})$ across all possible subsets $S\subseteq P\setminus\{k\}$. Although this suggests an exponential computational complexity, which becomes costly especially if $p$ is large, the second equality in Equation (26) reveals just linear complexity; for a proof we refer to Mayrhofer and Filzmoser, (2023). This allows for another insight into the Shapley value by comparing it to the squared Mahalanobis distance, which can be written as $\sum_{j,k=1}^{p}(x_{j}-\mu_{j})(x_{k}-\mu_{k})\omega_{jk}$. While the latter calculates an outlyingness measure by aggregating the contributions $(x_{j}-\mu_{j})(x_{k}-\mu_{k})\omega_{jk}$ of all variables for the entire observation, Equation (26) shows that a coordinate $\phi_{k}(x)$ of the Shapley value only considers the contributions that are associated with the $k$-th variable.
@@ 5.2 Shapley value for matrix-valued data
To define Shapley values for matrix-variate data, we can use the connection between the matrix and multivariate Mahalanobis distance; see Equation (6). Let $\bm{X}\in\mathbb{R}^{p\times q}$ be a matrix-variate sample with mean $\bm{\operatorname{M}}\in\mathbb{R}^{p\times q}$ and covariance matrices $\bm{\Sigma}^{\operatorname{row}}\in\operatorname{PSD}(p)$ and $\bm{\Sigma}^{\operatorname{col}}\in\operatorname{PSD}(q)$. The $pq$-dimensional vectorized observation is denoted as $\bm{x}=\operatorname{vec}(\bm{X})$, with mean $\bm{\mu}=\operatorname{vec}(\bm{\operatorname{M}})$ and covariance matrix $\bm{\Sigma}=\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}$. Based on Equation (26), we can obtain outlyingness contributions for every coordinate of $\bm{x}$ and hence for every cell of the matrix $\bm{X}$ by
@@
with $a=i+(l-1)p$ and $b=j+(k-1)p$. Analyzing the outlyingness scores for all $pq$ cells of an observation $\bm{X}\in\mathbb{R}^{p\times q}$ might be unpractical, especially when we want to compare multiple observations. Instead of the cellwise scores, we can consider row- and columnwise outlyingness scores for matrix-variate data. To obtain those scores, we could either sum up the cellwise contributions $\phi_{jk}(\bm{X})$, with $j=1,\ldots,p$ and $k=1,\ldots,q$, or replace an entire row $\bm{x}_{i.}$ of $\bm{X}$ with its corresponding mean, i.e., the $i$-th row of $\bm{\operatorname{M}}$. For the latter setting, we must adapt Equation (27) to the matrix case. The rowwise marginal contributions to the matrix Mahalanobis distance are defined as as
@@
$\begin{split}\phi_{j.}(\bm{X})&:=\sum_{S\subseteq P\setminus\{j\}}w(\left|S\right|)\Delta_{j}\operatorname{MMD}(\hat{\bm{X}}^{S})=\sum_{i=1}^{p}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{jl}-m_{jl})(x_{ik}-m_{ik})\omega_{ij}^{\operatorname{row}}\omega_{kl}^{\operatorname{col}},\end{split}$ (28)
Proof.
A proof for Equation (28) can be found in Supplement D. ∎
From the above, we conclude that adding up the contributions of all cells in a row yields identical scores as when we adapt the marginal contributions to a rowwise replacement. Using the first approach, we can sum up the columns and obtain columnwise outlying contributions. We can achieve the same by transposing the matrix in the latter case. The cellwise, rowwise, and columnwise Shapley values of all features can be computed efficiently using matrix operations as listed in Equations (29), (30), and (31), respectively. In Equation 29, the symbol $\circ$ refers to element-wise multiplication.
$\bm{\Phi}(\bm{X})$ $=(\bm{X}-\bm{\operatorname{M}})\circ\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})\bm{\Omega}^{\operatorname{col}}\in\mathbb{R}^{p\times q}$ (29)
$\bm{\phi}_{\operatorname{row}}(\bm{X})$ $=\operatorname{diag}(\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})\bm{\Omega}^{\operatorname{col}}(\bm{X}-\bm{\operatorname{M}})^{\prime})\in\mathbb{R}^{p}$ (30)
$\bm{\phi}_{\operatorname{col}}(\bm{X})$ $=\operatorname{diag}((\bm{X}-\bm{\operatorname{M}})^{\prime}\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})\bm{\Omega}^{\operatorname{col}})\in\mathbb{R}^{q}$ (31)
@@ 6 Simulations
In the simulation studies outlined in this section, our primary focus is to rigorously assess the performance of the MMCD estimators. We aim to validate their demonstrated theoretical properties and compare efficiency against ML estimators. Despite our initial intention to include various robust estimators as mentioned in Section 1, practical constraints arose as the relevant routines were exclusively accessible in Matlab, while our current framework operates within R. For the sake of consistency and practical implementation, we concentrate on comparing the efficiency of the raw and reweighted MMCD, alongside an in-depth analysis of MLE and (reweighted) MMCD on contaminated data. To ensure the highest possible breakdown point across all simulations and examples discussed in this paper, we set $h=\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$ for the MMCD estimators.
We generate samples from a matrix normal distribution with $(p,q)\in\{(5,20),(50,20)$, (100,50)}(100,50)\}( 100 , 50 ) }. The considered sample sizes are $n\in\{20,50,100,300,1000\}$. The mean matrix has all zero entries. For the rowwise covariance matrix we use randomly generated covariance matrices proposed by Agostinelli et al., (2015), $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}\in\operatorname{PSD}(p)$, generally yielding low correlations. For the columnwise covariance, we have $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)\in\operatorname{PSD}(q)$, with entries $\sigma_{jk}^{\operatorname{mix}}(0.7)=0.7^{\left|j-k\right|}$. Each simulation setting is replicated 100 times.
Using solely the clean samples, we assess the normal finite-sample efficiency by comparing the ratio
@@
where $D(\hat{\bm{\Sigma}}^{\operatorname{row}},\hat{\bm{\Sigma}}^{\operatorname{col}})$ denotes the Kullback-Leiber (KL) divergence of the estimators $\hat{\bm{\Sigma}}^{\operatorname{row}}$ and $\hat{\bm{\Sigma}}^{\operatorname{col}}$ in the matrix normal setting $\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$, and is given by
@@
with $\bm{\Omega}^{\operatorname{row}}=(\bm{\Sigma}^{\operatorname{row}})^{-1}$ and $\bm{\Omega}^{\operatorname{col}}=(\bm{\Sigma}^{\operatorname{col}})^{-1}$. Figure 2, shows that the efficiency of the raw MMCD estimators is below $0.5$ on average for both columnwise covariance matrices. In contrast, the reweighted estimators’ efficiency is above $0.5$ for $n=100$ and it rises to over $0.9$ as the sample size increases.
@@
For the setting with contamination, we generate two scenarios where 20% of the observations are outliers, with a mean matrix of either values of one or three, and with the same covariance matrices as for the regular observations. As evaluation measures, we use KL divergence (32) to analyze the quality of the covariance estimation. Additionally, we analyze outlier detection capabilities of the squared Mahalanobis distance based on MMCD and MLE covariance estimators, with the $\chi^{2}_{pq,0.99}$ quantile as a detection threshold. The performance is measured by the F-score, the harmonic mean of precision and recall. The results are presented in Figure 3, where the KL divergence is plotted on a log scale to enable a comparison between the robust MMCD estimators and the ML estimators.
@@
Figure 3: Comparison between F-Score and KL divergence of MMCD and ML estimators for simulated data from a matrix normal distribution with $20\%$ contamination.
In the upper row of Figure 3, the boxplots depict F-scores across various parameter configurations. Notably, the MMCD estimators exhibit improved performance as sample sizes increase across all settings, consistently outperforming ML estimators. However, for $(p,q)=(5,20)$, in a scenario involving a minor mean shift, the F-scores derived from MMCD exhibit some volatility with larger sample sizes. This situation arises due to the proximity of outliers to regular observations, posing challenges in their identification. Notably, a more pronounced mean shift significantly simplifies outlier detection. The KL divergence metric distinctly favors the robust estimators over classical ones. Yet, in the case of $(p,q)=(5,20)$ with a small mean shift and limited sample sizes, the median results between MMCD and ML estimators appear closer, albeit MMCD demonstrates significantly lower variability.
In Supplement E, we expand our analysis to cover additional simulation scenarios, incorporating diverse performance metrics such as precision, recall, Frobenius error, and the angle between eigenvalues of the covariance matrices.

### v2

@@ 5.1 Shapley values for multivariate data
  • Monotonicity: Let $\bm{\mu},\tilde{\bm{\mu}}\in\mathbb{R}^{p}$ be two vectors and $\bm{\Sigma},\tilde{\bm{\Sigma}}\in\operatorname{PDS}(p)$ be two matrices. If
@@
In words, the coordinate $\phi_{k}(\bm{x})$ of the Shapley value is the average marginal contribution of the $k$-th variable to the squared Mahalanobis distance and is obtained by averaging over all marginal outlyingness contributions $\Delta_{k}\operatorname{MD}^{2}(\hat{\bm{x}}^{S})$ across all possible subsets $S\subseteq P\setminus\{k\}$. Although this suggests an exponential computational complexity, which becomes costly, especially if $p$ is large, the second equality in Equation (24) reveals just linear complexity; for a proof we refer to Mayrhofer and Filzmoser, (2023). Equation (24) allows for another insight into the Shapley value by comparing it to the squared Mahalanobis distance, which can be written as $\sum_{j,k=1}^{p}(x_{j}-\mu_{j})(x_{k}-\mu_{k})\omega_{jk}$. While the latter calculates an outlyingness measure by aggregating the contributions $(x_{j}-\mu_{j})(x_{k}-\mu_{k})\omega_{jk}$ of all variables for the entire observation, Equation (24) shows that a coordinate $\phi_{k}(\bm{x})$ of the Shapley value only considers the contributions that are associated with the $k$-th variable.
@@ 5.2 Shapley value for matrix-valued data
To define Shapley values for matrix-variate data, we can use the connection between the matrix and multivariate Mahalanobis distance; see Equation (7). Let $\bm{X}\in\mathbb{R}^{p\times q}$ be a matrix-variate sample with mean $\bm{\operatorname{M}}\in\mathbb{R}^{p\times q}$ and covariance matrices $\bm{\Sigma}^{\operatorname{row}}\in\operatorname{PDS}(p)$ and $\bm{\Sigma}^{\operatorname{col}}\in\operatorname{PDS}(q)$. The $pq$-dimensional vectorized observation is denoted as $\bm{x}=\operatorname{vec}(\bm{X})$, with mean $\bm{\mu}=\operatorname{vec}(\bm{\operatorname{M}})$ and covariance matrix $\bm{\Sigma}=\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}$. Based on Equation (24), we can obtain outlyingness contributions for every coordinate of $\bm{x}$ and hence for every cell of the matrix $\bm{X}$ by
@@
with $a=i+(l-1)p$ and $b=j+(k-1)p$, and for $j=1,\ldots,p$ and $k=1,\ldots,q$. Using matrix operations, we can efficiently compute the $p\times q$ matrix containing the cellwise Shapley values $\phi_{jk}(\bm{X})$ as
$\bm{\Phi}(\bm{X})$ $=(\bm{X}-\bm{\operatorname{M}})\circ\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})\bm{\Omega}^{\operatorname{col}}\in\mathbb{R}^{p\times q},$ (26)
where $\circ$ refers to element-wise multiplication.
Next, we discuss how matrix affine transformations as in Equation (15) affect the cellwise Shapley values for matrix-variate data.
Proposition 5.2.1.
Let $\bm{X}\in\mathbb{R}^{p\times q}$ be a sample from $\mathcal{ME}(\bm{M},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}},g)$, $\bm{A}\in\mathbb{R}^{p\times p}$, $\bm{B}\in\mathbb{R}^{q\times q}$, $\bm{A},\bm{B}$ invertible, and $\bm{C}\in\mathbb{R}^{p\times q}$. Then, the cellwise Shapley values are not matrix affine equivariant, i.e., $\bm{\Phi}(\bm{A}\bm{X}\bm{B})\neq\bm{A}\bm{\Phi}\bm{X})\bm{B}$ for general positive definite $\bm{A}$ and $\bm{B}$. However, they are
  1. (a)
    shift invariant, i.e., $\bm{\Phi}(\bm{X}+\bm{C})=\bm{\Phi}(\bm{X})$,
  2. (b)
    scale invariant, i.e., if $\bm{A}$ and $\bm{B}$ are scaling matrices, thus diagonal matrices with non-zero entries, then $\bm{\Phi}(\bm{A}\bm{X}\bm{B})=\bm{\Phi}(\bm{X})$,
  3. (c)
    permutation equivariant, i.e., if $\bm{A}$ and $\bm{B}$ are permutation matrices, then $\bm{\Phi}(\bm{A}\bm{X}\bm{B})=\bm{A}\bm{\Phi}(\bm{X})\bm{B}$, and
The proofs are given in Supplement D. When considering gray-scale image data, shifting or rescaling the gray-scale information would not change the cellwise Shapley values. Further, exchanging rows and columns of the image; in particular mirroring or rotating the image by $90^{\circ}$, would equivalently transform the Shapley values. Similarly to the setting of cellwise outliers (Alqallaf et al.,, 2009), cellwise Shapley values are tied to the original coordinate system and are not matrix affine equivariant.
It can be preferable in some applications to obtain outlyingness explanations for a complete row or column of the matrix-valued observations, especially when we want to compare multiple observations. In the following, we show how Shapley values for rows can be obtained; Shapley values for columns can be computed based on the transposed matrix or by adapting the following notation accordingly for columns.
Consider again the set $P=\{1,\ldots,p\}$, and $S\subseteq P\setminus\{j\}$. The rowwise marginal contributions to the matrix Mahalanobis distance are defined as as
@@
$\phi_{j.}(\bm{X}):=$ $\sum_{S\subseteq P\setminus\{j\}}\frac{\left|S\right|!(p-\left|S\right|-1)!}{p!}\Delta_{j}\operatorname{MMD}(\hat{\bm{X}}^{S})$ (27)
$=$ $\sum_{i=1}^{p}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{jl}-m_{jl})(x_{ik}-m_{ik})\omega_{ij}^{\operatorname{row}}\omega_{kl}^{\operatorname{col}}=\sum_{k=1}^{q}\phi_{jk}(\bm{X}),$ (28)
A proof for Equation (28) can be found in Supplement D. Thus, a rowwise Shapley value is obtained by summing up the cellwise Shapley values for the corresponding row, which is equivalent to adapting the marginal contributions to a rowwise replacement. The vectors containing the rowwise or columnwise Shapley values can also be computed by
$\bm{\phi}_{\operatorname{row}}(\bm{X})$ $=\operatorname{diag}(\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})\bm{\Omega}^{\operatorname{col}}(\bm{X}-\bm{\operatorname{M}})^{\prime})\in\mathbb{R}^{p}\quad\text{and}$ (29)
$\bm{\phi}_{\operatorname{col}}(\bm{X})$ $=\operatorname{diag}((\bm{X}-\bm{\operatorname{M}})^{\prime}\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}})\bm{\Omega}^{\operatorname{col}})\in\mathbb{R}^{q},$ (30)
respectively. The properties listed in Proposition 5.2.1 also apply in this setting.
@@ 6 Simulations
In the simulation studies outlined in this section, our primary focus is to rigorously assess the performance of the MMCD estimators. We aim to validate their demonstrated theoretical properties and compare efficiency against ML estimators. Despite our initial intention to include various robust estimators as mentioned in Section 1, practical constraints arose as the relevant routines were exclusively accessible in Matlab, while our current framework operates within R. For the sake of consistency and practical implementation, we concentrate on comparing the efficiency of the raw and reweighted MMCD alongside an in-depth analysis of the MLEs, (reweighted) MMCD estimators, and MCD estimator based on the vectorized samples on contaminated data. To ensure the highest possible breakdown point across all simulations and examples discussed in this paper, we set $h=\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$ for the MMCD estimators and $h=\lfloor\nicefrac{{(n+pq+1)}}{{2}}\rfloor$ for the MCD estimator. We conduct 100 repetitions for each simulation setting and visualize the results through either line plots or boxplots. In the line plots, the solid lines represent average scores, while the shaded areas depict the one standard error regions.
Finite-sample efficiency.
To analyze the finite-sample efficiency, we generate samples from a centered matrix normal distribution with dimensions $(p,q)\in\{(5,20),(50,20)$, (100,50)}(100,50)\}( 100 , 50 ) } for various sample sizes $n\in\{20,50,100,300,1000\}$. For the rowwise covariance matrix we adopt the covariance matrices proposed by Agostinelli et al., (2015), denoted $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}\in\operatorname{PDS}(p)$, which have random entries and generally yield low correlations. For the columnwise covariance, we use $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)\in\operatorname{PDS}(q)$, with entries $\sigma_{jk}^{\operatorname{mix}}(0.7)=0.7^{\left|j-k\right|}$. We assess the normal finite-sample efficiency by comparing the ratio
@@
where $D(\hat{\bm{\Sigma}}^{\operatorname{row}},\hat{\bm{\Sigma}}^{\operatorname{col}})$ denotes the Kullback-Leiber (KL) divergence of the estimators $\hat{\bm{\Sigma}}^{\operatorname{row}}$ and $\hat{\bm{\Sigma}}^{\operatorname{col}}$ in the matrix normal setting $\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$, which is given by
@@
with $\bm{\Omega}^{\operatorname{row}}=(\bm{\Sigma}^{\operatorname{row}})^{-1}$ and $\bm{\Omega}^{\operatorname{col}}=(\bm{\Sigma}^{\operatorname{col}})^{-1}$. As shown in Figure 2, the efficiency of the raw MMCD estimators is below $0.5$ on average. In contrast, the reweighted estimators’ efficiency is above $0.5$ for $n=100$ and it rises to over $0.9$ as the sample size increases.
@@
Robustness and matrix size.
For the setting with contamination, we consider matrix-variate samples with $p\in\{2,\dots,30\}$ rows and $q=\{10,20,30\}$ columns for sample sizes $n\in\{100,1000\}$. The clean data are generated from a centered matrix normal distribution with $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}$ and $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)$. A fraction, $\varepsilon=0.1$, of the clean data is replaced by outliers, sampled from a matrix normal distribution with a mean matrix where all entries are equal to $\gamma=1$. The covariance matrices of the outliers are the same as for the regular observations.
We use KL divergence (31) to analyze the quality of the covariance estimation. Additionally, we analyze outlier detection capabilities of the squared Mahalanobis distance based on the estimators, with the $\chi^{2}_{pq,0.99}$ quantile as a detection threshold. We also include the Mahalanobis distances based on true parameters used to generate the data as a benchmark and measure performance by precision and recall. Due to the excessively long computation times of the Fast-MCD procedure in higher-dimensional scenarios, we used the deterministic MCD (Hubert et al.,, 2012) when $pq>300$. Since the MCD estimator requires $n>pq$, it is only computed for those settings.
@@
Figure 3: Comparison of precision, recall, and KL divergence for ML and MMCD estimators, (deterministic) MCD estimators with vectorized data, and true parameters as a benchmark for outlier detection for simulated data from a matrix normal distribution with $10\%$ contamination.
Figure 3 shows that the MMCD estimators have lower KL divergence than the competing methods and attain a recall similar to the benchmark approach based on the true parameters used to generate the data across all settings. The precision of the MMCD estimators depends on the dimensionality of the matrix-variate samples as well as on the sample size. For $n=100$, the precision decreases with increasing dimensionality $pq$, but the effect is mitigated by an improving performance when $\max\{\nicefrac{{p}}{{q}},\nicefrac{{q}}{{p}}\}$ is small. For $n=1000$, the precision is close to the precision based on the true parameters. This suggests that for small sample sizes, a correction similar to the one proposed by Pison et al., (2002) for the MCD could lead to a better performance. In the matrix-variate setting, such a correction would not only be dependent on $pq$ and $n$ but also on $\nicefrac{{p}}{{q}}$ and $\nicefrac{{q}}{{p}}$.
For small $p$ and $q$, the comparison between the MMCD estimators and the MCD for the vectorized observations is of special interest. For $n=1000$ and $q=10$ they have a similar recall when $p\leq 6$, and for $q\in\{20,30\}$ the MCD estimators show substantial improvements when the deterministic MCD approach is used instead of the Fast-MCD. This can be explained by the dependence of the Fast-MCD on the robustness of the initial solutions, and with an increasing $pq$, the probability of obtaining a clean subset becomes very small. The MCD estimator shows a steep drop in precision as $p$ increases when Fast-MCD is used. For the deterministic MCD, we see a trade-off between precision and recall with increasing dimensionality, but the KL divergence remains high. With increasing dimensionality, even the nonrobust matrix MLEs outperform the MCD estimator which highlights the importance of respecting the inherent data structure of matrix-variate observations.
Robustness and contamination type.
In addition to the shift outliers we also consider block and cell contamination for matrix normal samples of size $(p,q)=(5,20)$. In all three settings, we consider a fraction of $\varepsilon=0.1$ contaminated samples. Let $\bm{X}=(x_{jk}),j=1,\dots,5,k=1,\dots,20$, denote a sample from a centered matrix normal distribution with rowwise covariance $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}$ and columnwise covariance $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)$. For block contamination, we replaced the top left $2\times 5$ block, corresponding to the entries $x_{jk},j=1,2,k=1,\dots,5$, with entries from a shifted matrix normal distribution with a mean matrix where all entries are equal to $\gamma=1$ and covariance matrices corresponding to the top left block of $\bm{\Sigma}^{\operatorname{row}}$ and $\bm{\Sigma}^{\operatorname{col}}$. For cell contamination, a fraction of $0.1$ of the cells of the outlying observations are randomly permuted. The shift outliers are generated with a mean shift $\gamma=1$ as before.
Figure 4: Precicion, recall, and logarithm of KL divergence comparing block, cell, and sample contamination.
Figure 4 shows that the MMCD estimators are better suited for outlier detection and yield more robust covariance estimates than the matrix MLEs as well as the MCD estimator on the vectorized observations. Overall, the results are similar across all three simulation scenarios, only for mean shift contamination we see higher variation than in the other two settings. This is likely because the block and cell contamination interfere with covariance estimation more profoundly, i.e., the KL divergence of the matrix MLEs is highest for block contamination followed by cell and shift contamination.
The supplementary materials E provide in-depth simulation studies that expand upon the scenarios discussed in this section. These simulations analyze the effects of the level of contamination and mean shifts for multiple types of covariance matrices. Additionally, we extend our analysis beyond the normal model to include samples generated from a matrix t-distribution, examining performance across a range of degrees of freedom. For this scenario, we also compute the ML estimators for the matrix t-distribution (Thompson et al.,, 2020). We include a summary of computation time and consider additional performance metrics, such as the F-score (harmonic mean of precision and recall), Frobenius error, and the angle between eigenvalues of covariance matrices.